# Muddit: Liberating Generation Beyond Text-to-Image with a Unified Discrete Diffusion Model

### v1

@@ Abstract
1Peking University, 2TeleAI, China Telecom, 3National University of Singapore, 4Princeton University🖂:  jinbin.bai@u.nus.edu
@@ 1 Introduction
Recent work like Dual-Diffusion [29] (Fig. 1(c)) claims to unify modalities under discrete diffusion, but it ultimately relies on continuous diffusion for image generation via Stable Diffusion 3, a continuous diffusion paradigm. This fundamental mismatch in generative principles undermines its claim of true unification. UniDisc [48](Fig. 1(d)), takes a more promising step by applying discrete diffusion [^1] MaskGIT, MaskAR, RandomAR, and Discrete Diffusion share significant conceptual and practical overlaps, often differing only in decoding order or architectural nuances. We elaborate on their connections in the next section. While Meissonic [5] follows the naming convention of MaskGIT [8], we standardize terminology in this paper by referring to all such models under the umbrella of Discrete Diffusion. over unified token spaces. This allows parallel refinement of text and image tokens, improving inference efficiency and enabling more flexible conditioning. However, the overall generation quality of UniDisc remains far from satisfactory. For example, it struggles to produce high-resolution $1024\times 1024$ images, fails to match the fidelity of early diffusion models such as Stable Diffusion 1.5, and lacks support for vision-language reasoning tasks such as visual question answering (VQA). These limitations expose the second “dark cloud”: the absence of strong pre-trained discrete diffusion backbone models: Unlike established unified autoregressive models that leverage powerful pretrained large language models, current unified discrete diffusion models are typically trained from scratch on mixed-modality tokens, which limits both their generative fidelity and transferability. Without modular components carrying rich pixel-level priors, these models face generalization and scalability bottlenecks.
@@ 3.1 Discrete Diffusion with Unified Image and Text Perspective
In discrete diffusion, a sample $x\in\mathcal{X}$ is treated as a one-hot vector $\mathbf{x}$, where $\mathcal{X}=\{1,\dots,N\}$. For language models, $N$ equals the vocabulary size. While for image models, $N$ is the number of discrete image–token IDs obtained from a tokenizer or VQ-codebook. At each diffusion step, we stochastically corrupt the tokens, gradually transforming the data distribution into a maximally entropic categorical prior; the generative model then learns to invert this corruption. Following recent works [36, 5] that cast token corruption as a continuous–time Markov chain (CTMC) over the finite alphabet $\mathcal{X}$, we let
@@
Reverse process. For any $10<s<t<$ 1, the CTMC induces an analytic posterior
@@
Training Objective. We employ a masked-token predictor $x_{\theta}(x_{t},\alpha_{t})\!\approx\!\mathbf{x}$, which leads to the continuous-time negative ELBO
[…]
@@
During generation, we start from an all-mask sequence ($t=1$) and integrate the reverse CTMC towards $t=0$, repeatedly replacing every masked position with the model’s categorical prediction. Because the corruption schedule and objective are identical for any discrete alphabet $\mathcal{X}$, the same diffusion backbone unifies text and image generation. In the following section, we present Muddit, a unified framework that leverages discrete diffusion to model the generation tasks for both text and image jointly.
@@ 3.2.1 Unified Architecture
Figure 2: The training and inference architecture of Muddit. (a) During training, we randomly mask tokens from one of the two modalities. MM-DiT is trained to predict the masked tokens using a re-weighted cross-entropy loss, which jointly optimizes both the MM-DiT backbone and a lightweight text decoder. (b) In text-to-image inference, we initialize the image latent features using all-masked tokens and iteratively predict each latent token via MM-DiT. (c) In image-to-text inference, we similarly initialize all text tokens as masked and generate the text through the same iterative decoding process. Specifically for VQA tasks, we append mask token IDs to the end of the question and predict all masked token IDs as the final answer.
@@ 3.2.2 Unified Training
    $\gamma_{t}$ must be a continuous function, bounded between 00 and $1$, for $t\in[0,1]$.
@@
[…]
where all symbols are as in Eq. 4 but the $\mathtt{G}$ now receives the cross-modal condition $\mathbf{c}$ as an additional input. Key point: switching from text \!$\to$ image to image \!$\to$ text merely changes the conditioning signal $\mathbf{c}$; the loss Eq. 6 itself is unchanged. This symmetry keeps optimization identical across tasks and allows us to train a single parameter set jointly for both generation directions. During inference we again start from an all-mask sequence ($t{=}1$) and integrate the reverse CTMC towards $t{=}0$, feeding in the desired condition $\mathbf{c}$ to obtain either an image or a sentence from the same diffusion backbone.
@@ 3.2.3 Unified Inference
(i) Text $\boldsymbol{\to}$ image. Given a text prompt $\mathtt{tp}\!\in\!\mathcal{T}$, the text encoder $\mathtt{E_{txt}}$ produces a text token embedding $\mathtt{c_{txt}}=\mathtt{E_{txt}}(\mathtt{tp})$. Starting from a fully masked sequence $x_{1}$, the generator produces logits
@@
(ii) Image $\boldsymbol{\to}$ text. For captioning, an input image $I\!\in\!\mathcal{I}$ is tokenized by the image encoder $\mathtt{E_{img}}$: $\mathtt{c_{img}}=\mathtt{E_{img}}(I)$. The generator now conditions on the visual tokens while progressively decoding text:
@@
(iii) Image + question $\boldsymbol{\to}$ answer (VQA). For visual–question answering we supply both an image and a question: $\mathtt{c_{img}}=\mathtt{E_{img}}(I)$ and $\mathtt{c_{txt}}=\mathtt{E_{txt}}(q)$. They are concatenated and fed to the generator, which outputs logits over answer tokens $x_{k}$:
@@ 4.2 Text-to-Image Generation
Model Text Gen Arch Image Gen Arch Params (B) Overall $\uparrow$ Objects $\uparrow$ Counting $\uparrow$ Colors $\uparrow$ Position $\uparrow$ Color $\uparrow$ Attribution
@@ 4.3 Image-to-Text Generation
We present a comprehensive comparison of our model Muddit against other multimodal models across four benchmarks: MS-COCO (image captioning) [31], VQAv2 [2], MME [19], and GQA [24] in Tab. 2. Notably, Muddit is the first unified model to employ discrete diffusion for both image and text generation, demonstrating that this approach is not only viable but also highly competitive.
@@
Model Params Text Gen Image Gen MS-COCO VQAv2 MME GQA
(B) Arch Arch CIDEr $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$
InternVL-2.0 [14] 8 AR - - - 1648.1 61.0
LLaVA-Next [33] 13 AR - - 82.8 1575.0 65.4
BLIP-2 [27] 13 AR - - 65.0 1293.8 41.0
QWEN-VL [4] 7 AR - - 78.2 1487.5 57.5
OpenFlamingo [3] 9 AR - 65.5 43.5 - -
Flamingo [1] 9 AR - 79.4 51.8 - -
Chameleon [50] 7 AR AR 18.0 - - -
LWM [34] 7 AR AR - 55.8 - 44.8
Show-O (256$\times$256) [56] 1.3 AR Discrete Diff. - 64.7 1014.9 54.2
Show-O (512$\times$512) [56] 1.3 AR Discrete Diff. - 69.4 1097.2 58.0
Transfusion [60] 7 AR Diffusion 29.0 - - -
D-DiT (256$\times$256) [29] 2 Discrete Diff. Diffusion - 59.5 897.5 55.1
D-DiT (512$\times$512) [29] 2 Discrete Diff. Diffusion 56.2 60.1 1124.7 59.2
UniDisc [49] 0.33 Discrete Diff. Discrete Diff. 46.8 - - -
Muddit (512$\times$512) 1 Discrete Diff. Discrete Diff. 59.7 67.7 1104.6 57.1
Table 2: Evaluation of image-to-text generation and visual question answering.
@@ 4.4 Ablation Study and Analysis
[…]
@@
MS-COCO CIDEr $\uparrow$ 50.2 51.2 58.4 58.3
VQAv2 $\uparrow$ 62.1 65.8 67.8 67.9
@@
Metric w/o joint training w/ joint training
GenEval $\uparrow$ 28.3 61.6
MS-COCO CIDEr $\uparrow$ 59.4 58.4
VQAv2 $\uparrow$ 69.2 67.8
@@
GenEval $\uparrow$ 51.4 58.1 59.2 61.6 61.5 61.4 60.8
MS-COCO CIDEr $\uparrow$ 43.4 58.5 58.6 58.4 59.2 60.0 59.7
VQAv2 $\uparrow$ 68.3 68.4 68.5 67.8 67.5 67.6 67.7
@@ 4.5 Inference Time Analysis
As shown in Fig. 6, autoregressive multimodal models are inherently limited by token-by-token decoding, which constrains their inference speed. Muddit overcomes this bottleneck with a parallel discrete diffusion decoder, reducing average latency to just 1.49 seconds, achieving a 4$\times$ to 11$\times$ speedup over competitive baselines (4.2$\times$ faster than Qwen-2.5-VL, 5.6$\times$ than Show-o, 8.1$\times$ than BLIP-2, and 10.9$\times$ than LLaVA-1.6).

### v3

@@ Abstract
1Peking University, 2TeleAI, 3National University of Singapore, 4Princeton University🖂:  jinbin.bai@u.nus.edu
@@ 1 Introduction
Recent work like Dual-Diffusion [29] and Diffuse Everything [47] (Fig. 1(c)) claims to unify modalities under discrete diffusion, but they ultimately rely on continuous diffusion for image generation. This fundamental mismatch in generative principles undermines their claim of true unification. UniDisc [53](Fig. 1(d)), takes a more promising step by applying discrete diffusion [^1] MaskGIT, MaskAR, RandomAR, and Discrete Diffusion share significant conceptual and practical overlaps, often differing only in decoding order or architectural nuances. We elaborate on their connections in the next section. While Meissonic [5] follows the naming convention of MaskGIT [8], we standardize terminology in this paper by referring to all such models under the umbrella of Discrete Diffusion. over unified token spaces. This allows parallel refinement of text and image tokens, improving inference efficiency and enabling more flexible conditioning. However, the overall generation quality of UniDisc remains far from satisfactory. For example, it struggles to produce high-resolution $1024\times 1024$ images, fails to match the fidelity of early diffusion models such as Stable Diffusion 1.5, and lacks support for vision-language reasoning tasks such as visual question answering (VQA). These limitations expose the second “dark cloud”: the absence of strong pre-trained discrete diffusion backbone models: Unlike established unified autoregressive models that leverage powerful pretrained large language models, current unified discrete diffusion models are typically trained from scratch on mixed-modality tokens, which limits both their generative fidelity and transferability. Without modular components carrying rich pixel-level priors, these models face generalization and scalability bottlenecks.
@@ 3.1 Discrete Diffusion with Unified Image and Text Perspective
In discrete diffusion, a sample $x\in\mathcal{X}$ is treated as a one-hot vector $\mathbf{x}$, where $\mathcal{X}=\{1,\dots,N\}$. For language models, $N$ equals the vocabulary size. While for image models, $N$ is the number of discrete image–token IDs obtained from a tokenizer or VQ‑codebook. At each diffusion step, we stochastically corrupt the tokens, gradually transforming the data distribution into a maximally entropic categorical prior; the generative model then learns to invert this corruption. Following recent works [37, 5] that cast token corruption as a continuous–time Markov chain (CTMC) over the finite alphabet $\mathcal{X}$, we let
@@
Reverse process. For any $0<s<t<1$, the CTMC induces an analytic posterior
@@
Training Objective. We employ a masked‐token predictor $x_{\theta}(x_{t},\alpha_{t})\!\approx\!\mathbf{x}$, which leads to the continuous‐time negative ELBO [42, 49, 50]
$\mathcal{L}_{\mathrm{NELBO}}=\mathbb{E}_{q(x_{t}\mid\mathbf{x})}\;\Bigl[\int_{0}^{1}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\,\log\bigl(x_{\theta}(x_{t},\alpha_{t})\!\cdot\!\mathbf{x}\bigr)\,dt\Bigr],$ (4)
@@
During generation, we start from an all‐mask sequence ($t=1$) and integrate the reverse CTMC towards $t=0$, repeatedly replacing every masked position with the model’s categorical prediction. Because the corruption schedule and objective are identical for any discrete alphabet $\mathcal{X}$, the same diffusion backbone unifies text and image generation. In the following section, we present Muddit, a unified framework that leverages discrete diffusion to model the generation tasks for both text and image jointly.
@@ 3.2.1 Unified Architecture
Figure 2: The training and inference architecture of Muddit. (a) During training, we randomly mask tokens from one of the two modalities. $M$-DiT is trained to predict the masked tokens using a re-weighted cross-entropy loss, which jointly optimizes both the MM-DiT backbone and a lightweight text decoder. (b) In text-to-image inference, we initialize the image latent features using all-masked tokens and iteratively predict each latent token via $M$-DiT. (c) In image-to-text inference, we similarly initialize all text tokens as masked and generate the text through the same iterative decoding process. Specifically for VQA tasks, we append mask token IDs to the end of the question and predict all masked token IDs as the final answer.
@@ 3.2.2 Unified Training
    $\gamma_{t}$ must be a continuous function, bounded between 0 and $1$, for $t\in[0,1]$.
@@
$\mathcal{L}_{\mathrm{unified}}=\mathbb{E}_{q(x_{t}\mid\mathbf{x})}\;\Bigl[\int_{0}^{1}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\,\log\!\bigl(\mathtt{G}(x_{t},\alpha_{t},\mathbf{c})\!\cdot\!\mathbf{x}\bigr)\,dt\Bigr],$ (6)
where all symbols are as in Eq. 4 but the $\mathtt{G}$ now receives the cross-modal condition $\mathbf{c}$ as an additional input. Key point: switching from text $\to$ image to image $\to$ text merely changes the conditioning signal $\mathbf{c}$; the loss Eq. 6 itself is unchanged. This symmetry keeps optimization identical across tasks and allows us to train a single parameter set jointly for both generation directions. During inference we again start from an all-mask sequence (t=1t{=}1) and integrate the reverse CTMC towards t=0t{=}0, feeding in the desired condition $\mathbf{c}$ to obtain either an image or a sentence from the same diffusion backbone.
@@ 3.2.3 Unified Inference
[…]
@@
[…]
@@
[…]
@@ 4.2 Text-to-Image Generation
Model
Text Gen
Arch
Image Gen
Arch
Params
Overall $\uparrow$ Objects $\uparrow$ Counting $\uparrow$ Colors $\uparrow$ Position $\uparrow$
Color $\uparrow$
Attribution
@@ 4.3 Image-to-Text Generation
We present a comprehensive comparison of our model Muddit against other multimodal models across six benchmarks: MS-COCO (image captioning) [31], VQAv2 [2], MME [19], MMBench [36], GQA [24], and MMMU [63] in Tab. 2. Notably, Muddit is the first unified model to employ discrete diffusion for both image and text generation, demonstrating that this approach is not only viable but also highly competitive.
@@
Model Params (B) Text Gen Arch Image Gen Arch MS-COCO CIDEr $\uparrow$ VQAv2 Acc. $\uparrow$ MME Acc. $\uparrow$ MMB Acc. GQA Acc. $\uparrow$ MMMU Acc.
InternVL-2.0 [14] 8 AR - - - 1648.1 81.7 61.0 49.3
LLaVA-Next [33] 13 AR - - 82.8 1575.0 70.0 65.4 36.2
BLIP-2 [27] 13 AR - - 65.0 1293.8 - 41.0 34.4
QWEN-VL [4] 7 AR - - 78.2 1487.5 - 57.5 35.9
OpenFlamingo [3] 9 AR - 65.5 43.5 - - - 28.7
Flamingo [1] 9 AR - 79.4 51.8 - - - -
Chameleon [55] 7 AR AR 18.0 - - 19.8 - -
LWM [34] 7 AR AR - 55.8 - - - -
Show-O ($\times256$) [61] 1.3 AR Discrete Diff. - 64.7 1014.9 - 54.2 -
Show-O ($\times512$) [61] 1.3 AR Discrete Diff. - 69.4 1097.2 - 58.0 27.4
Transfusion [66] 7 AR Diffusion 29.0 - - - - -
D-DiT ($\times256$) [29] 2 Discrete Diff. Diffusion - 59.5 897.5 - 55.1 -
D-DiT ($\times512$) [29] 2 Discrete Diff. Diffusion 56.2 60.1 1124.7 - 59.2 -
UniDisc [54] 1.4 Discrete Diff. Discrete Diff. 46.8 - - - - -
Muddit ($\times512$) 1 Discrete Diff. Discrete Diff. 59.9 68.2 1107.4 58.4 57.5 27.6
Table 2: Evaluation of image captioning, visual question answering on multimodal benchmarks.
@@ 4.4 Ablation Study and Analysis
Analysis of the inference timesteps. As shown in Tab. 5, increasing the number of diffusion steps generally improves performance, with most metrics plateauing around $T=32\text{--}50$. In particular, GenEval and CIDEr scores improve substantially from $T=8$ to $T=32$, though the marginal gains diminish thereafter. VQAv2 remains largely stable across timesteps, suggesting that fewer steps suffice for discriminative tasks. Overall, a moderate number of steps offers a favorable trade-off between accuracy and efficiency.
@@
MS-COCO CIDEr $\uparrow$ 51.4 52.1 59.9 58.8
VQAv2 $\uparrow$ 62.7 66.2 68.2 68.4
@@
Metric T2I only I2T only Joint training
GenEval $\uparrow$ 59.3 28.3 61.6
MS-COCO $\uparrow$ - 60.1 59.9
VQAv2 $\uparrow$ - 69.1 68.2
@@
GenEval $\uparrow$ 51.6 58.5 59.3 61.9 61.7 61.5 61.1
MS-COCO CIDEr $\uparrow$ 43.6 59.3 59.4 59.7 60.1 60.4 59.9
VQAv2 $\uparrow$ 53.9 57.4 62.3 65.4 66.8 67.7 68.2
@@ 4.5 Inference Time Analysis
[…]